# High-probability Convergence Guarantees of Decentralized SGD

### v3

@@ 4 Proof Outlines and Discussion
We use Lemma 3 to control the last term on the RHS, while to deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ and remove the need for bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [37, 39, 2], i.e., we subtract $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of the inner product is absorbed by the left-hand side. The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix E for details.
@@
We use (4), Lemma 4, and a careful analysis to show that the MGF of $F^{t}$ satisfies the conditions of Lemma 5, which we can then apply. The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix F for details.
@@
losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix I. On the other hand, Lemma 5 provides a much sharper bound on the MGF, ensuring that linear speed-up is achieved. Similar observations hold with respect to [41], see Appendix I for details.
@@
Our bounds contain a factor of $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for MSE or HP bounds in [33, 44]. This stems from Lemma 2, where $d$ appears in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4). One way to relax this dependence to $\log(d)$ is to assume that the noise is norm-sub-Gaussian, which comes with the cost of a slightly stronger noise condition, see, e.g., Corollary 7 in [27]; see also Appendix K for a further discussion on the dimension dependence.
@@ Appendix A Introduction
The Appendix contains results omitted from the main body. Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, Section G provides numerical results, Section H derives the transient times, Section I provides a discussion on the novel results needed for achieving linear speed-up, Section J discusses the extension to time-varying networks, while Section K contains further discussion on dimension dependence in our bounds.
Appendix B Useful Inequalities
@@
Finally, we complete this section with the following important consequences of (A6), see, e.g., [55].
@@
Appendix C Technical Results
@@
    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
@@
    Combining the previous result with Lemma 1 from [27], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
    $\mathbb{P}\big(\|\overline{z}^{t}\|>\epsilon\big)\leq 2\exp\bigg(-\frac{n\epsilon^{2}}{24\sigma^{2}d}\bigg).$
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., Proposition 2.5.2 in [67], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
@@
where the second inequality follows from the fact that $b_{t}\in(0,1)$ and Proposition 1. Unrolling the recursion (C) and noting that $Y_{1}=X_{1}$, since $X_{1}>0$ is deterministic, we get
@@
Plugging (13) into (C) completes the proof. ∎
Appendix D Analysis Setup
@@
Appendix E Proofs for Non-convex Costs
@@
where we used Proposition 1 and (A3) in the last inequality. Plugging (20) in (E) gives the desired result. ∎
@@
Theorem 3.4.
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
@@
$=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right]$
$\leq\mathbb{E}\left[\exp(B_{2,T-1})\left(\mathbb{E}\left[\exp\left(\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right)^{\nicefrac{{360\alpha_{T}^{2}\sigma^{2}dL}}{{n}}}\right],$
where the last inequality follows by choosing $\alpha_{t}\leq\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$ and using Proposition 1. Using Lemma 2, we get
$\mathbb{E}[\exp(B_{2,T})]\leq\mathbb{E}\left[\exp(B_{2,T-1})\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\right]=\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\mathbb{E}[\exp(B_{2,T-1})].$
Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{3\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}\right)$. Finally, to bound $\mathbb{E}[\exp(B_{3,T})]$, we can proceed in the same way, using the conditional independence of noise across agents (recall (A4)) to note that, if $\alpha_{t}\leq\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2/3}L^{2/3}\sqrt[3]{9}}$, then $\mathbb{E}[\exp(B_{3,T})]\leq\exp\left(\frac{9\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\right)$. Combining everything, we get
$\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg).$ (29)
@@
$\mathbb{P}\left(M_{T}>\epsilon\right)\leq\exp(-\epsilon)\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(-\epsilon+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg),$
@@
$M_{T}\leq\log(\nicefrac{{1}}{{\delta}})+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
Recalling the definition of $M_{T}$ and using the fact that the sequence of step-sizes in non-increasing we then have, with probability at least $1-\delta$
@@
$+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
@@
$+\frac{1920\sigma^{2}dL}{n\alpha_{T}T}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{48\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\alpha_{T}T(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
@@
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\alpha\sigma^{2}dL}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dLT}}$, then with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dLT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dLT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{1}{CT}$. Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120\big)}{\sqrt{nT}}$
    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$

### v5

@@ 4 Proof Outlines and Discussion
We use Lemma 3 to control the last term on the RHS, while to deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ and remove the need for bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [40, 42, 2], i.e., we subtract $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of the inner product is absorbed by the left-hand side. The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix F for details.
@@
We use (4), Lemma 4, and a careful analysis to show that the MGF of $F^{t}$ satisfies the conditions of Lemma 5, which we can then apply. The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix G for details.
@@
losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix J. On the other hand, Lemma 5 provides a much sharper bound on the MGF, ensuring that linear speed-up is achieved. Similar observations hold with respect to [44], see Appendix J for details.
@@
Our bounds contain a factor of $\log(d)$, introducing a mild dependence on the problem dimension, which is not the case for either MSE results [35], or HP bounds in [47]. This dependence stems from Lemma 2, where we show that variance-reduction is achieved in the HP sense, with $\log(d)$ appearing in the exponent bounding the MGF of $\overline{z}^{t}$. While mild, it is not clear if this dependence on the problem dimension can be completely removed for sub-Gaussian vectors, as noted also in [29]. For a further discussion on sub-Gaussian vectors and dimension dependence, see Appendix L.
@@ Appendix A Introduction
Appendix contains results omitted from the main body. Section B discusses the differences between MSE and HP guarantees, Section C collects some important facts used in our proofs, Section D provides some technical results, Section E defines some notions used in the analysis, Sections F and G provide proofs for non-convex and strongly convex costs, respectively, Section H provides numerical results, Section I derives the transient times, Section J provides a discussion on the novel results needed for achieving linear speed-up, Section K discusses the extension to time-varying networks, while Section L contains further discussion on dimension dependence in our bounds and the notion of sub-Gaussianity used in our work.
Appendix B On MSE and HP Results
Although both HP and MSE results establish convergence, the nature of these guarantees is very different. In particular, MSE results quantify the average behaviour across many runs of an algorithm, while HP results quantify the behaviour of an individual run. As mentioned briefly in the introduction, this distinction is very important in huge-scale applications like LLM training, where it is often impossible to perform more than a single training run, both resource- and time-wise. For example, [54] estimate that the GPT-3 model with 175 billion parameters was trained for 34 days, while the variant with 1 trillion parameters can be trained for approximately 84 days. This is further exacerbated by phenomena like heavy-tailed noise, frequently observed during training of deep learning models and transformers, e.g., [89, 72], which can cause the performance of an individual run to significantly deviate from the average performance. Hence, strong guarantees with respect to an individual run are very important in such applications. While MSE results can be used to provide guarantees with respect to an individual run, they can be quite loose compared to HP bounds. In particular, if we have a MSE bound of the form $\mathbb{E}[X_{t}^{2}]=1/t$, Chebyshev’s inequality implies $\mathbb{P}\big(X_{t}^{2}>\nicefrac{{1}}{{\delta t}}\big)\leq\delta$, which shows an inversely proportional dependence on the confidence level $\delta\in(0,1)$. On the other hand, HP results establish guarantees of the type $\mathbb{P}(X_{t}^{2}>\log(\nicefrac{{1}}{{\delta}})/t)\leq\delta$, having a much milder, logarithmic dependence on $1/\delta$. Equivalently, for any threshold $\epsilon>0$, MSE + Chebyshev implies $\mathbb{P}(X_{t}^{2}>\epsilon)\leq\nicefrac{{1}}{{t\epsilon}}$, while HP gives $\mathbb{P}(X_{t}^{2}>\epsilon)\leq\exp(-t\epsilon)$, providing a much sharper bound on the tail probability. Since our goal is to show that $X_{t}^{2}\leq\epsilon$ (e.g., if $X_{t}=\|x^{t}-x^{\star}\|$, or if $X_{t}=\min_{k\in[t]}\|\nabla f(x^{k})\|$), then $\mathbb{P}(X_{t}^{2}>\epsilon)$ represents the probability of failure, with MSE + Chebyshev implying that the probability of failure decays polynomially in $t$, while HP bound implies that the probability of failure decays exponentially fast in $t$, giving much stronger guarantees on the convergence of an individual run.
Appendix C Useful Inequalities
@@
Next, we state some important consequences of (A6), see, e.g., [59].
@@
Finally, we provide a useful martingale concentration result from [29].
Proposition 9.
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{d}$ be random vectors and denote by $\mathcal{F}_{n}=\sigma\big(\{X_{1},\ldots,X_{n}\}\big)$ the corresponding filtration. If for all $i\in[n]$, the vector $X_{i}$ satisfies $\mathbb{E}[X_{i}\>|\>\mathcal{F}_{i-1}]=0$ and
$\mathbb{P}(\|X_{i}\|>\epsilon\>|\>\mathcal{F}_{i-1})\leq 2\exp\bigg(-\frac{t^{2}}{2\sigma_{i}^{2}}\bigg),$
for some $\mathcal{F}_{i-1}$-measurable $\sigma_{i}>0$ and all $\epsilon>0$, then for any $\delta\in(0,1)$, we have
$\mathbb{P}\Big(\|\sum_{i\in[n]}X_{i}\|^{2}>2\sqrt{2}\sum_{i\in[n]}\sigma_{i}^{2}\log(\nicefrac{{2d}}{{\delta}})\Big)\leq\delta.$
Appendix D Technical Results
@@
    The average noise $\overline{z}^{t}$ satisfies $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{15\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq 2d\exp(1)$.
@@
    From (A4) we know that conditioned on $\mathcal{F}_{t}$, the vectors $z_{1}^{t},\ldots,z_{n}^{t}$ are independent and zero-mean. Moreover, for each $i\in[n]$ and any $\epsilon>0$, we have
    $\mathbb{P}\big(\|z^{t}_{i}\|>$ $\epsilon\>|\>\mathcal{F}_{t}\big)=\mathbb{P}\bigg(\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{2\sigma_{i}^{2}}\bigg)>\exp\bigg(\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg)\stackrel{{\scriptstyle(a)}}{{\leq}}\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{2\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg]$
    $\stackrel{{\scriptstyle(b)}}{{\leq}}\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/2}\stackrel{{\scriptstyle(c)}}{{\leq}}\exp\bigg(\frac{1}{2}-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\stackrel{{\scriptstyle(d)}}{{\leq}}2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg),$
    where $(a)$ follows from Markov’s inequality, $(b)$ follows from Proposition 1, in $(c)$ we use (A4), while $(d)$ follows form the fact that $\exp(1/2)\leq 2$. Therefore, vectors $z_{1}^{t},\ldots,z_{n}^{t}$ satisfy the conditions of Proposition 9 with respect to $\mathcal{F}_{t}$, and it follows that, for any $\delta\in(0,1)$
    $\mathbb{P}\bigg(\|\overline{z}^{t}\|^{2}>\frac{2\sqrt{2}\sigma^{2}\log(\nicefrac{{2d}}{{\delta}})}{n}\bigg)\leq\delta,$
    or equivalently, for any $\epsilon>0$, we have
    $\mathbb{P}(\|\overline{z}^{t}\|^{2}>\epsilon)\leq 2d\exp\bigg(-\frac{n\epsilon}{2\sigma^{2}\sqrt{2}}\bigg)\leq 2d\exp\bigg(-\frac{n\epsilon}{3\sigma^{2}}\bigg).$ (8)
    Next, define $Y=\frac{\sqrt{n}\|\overline{z}^{t}\|}{\sigma\sqrt{3}}$. Using the layer-cake formula for the expectation of a non-negative random variable, we then have, for any integer $p\geq 1$
    $\mathbb{E}\big[\|Y\|^{2p}\big]$ $=\int_{0}^{\infty}\mathbb{P}\big(\|Y\|^{2p}>y\big)dy\stackrel{{\scriptstyle(a)}}{{=}}p\int_{0}^{\infty}s^{p-1}\mathbb{P}\big(\|Y\|^{2}>s\big)ds$
    $=p\int_{0}^{\infty}s^{p-1}\mathbb{P}\bigg(\|\overline{z}^{t}\|^{2}>\frac{3\sigma^{2}s}{n}\bigg)ds\stackrel{{\scriptstyle(b)}}{{\leq}}2dp\int_{0}^{\infty}s^{p-1}\exp(-s)ds\stackrel{{\scriptstyle(c)}}{{\leq}}2dp^{p+1},$ (9)
    where $(a)$ follows by introducing the substitution $y=s^{p}$, $(b)$ follows from (8), while $(c)$ follows from the definition of $\Gamma$ function and the fact that $\Gamma(p)\leq p^{p}$. For any $c>0$, we then have
    $\mathbb{E}\bigg[\exp\bigg(\frac{cn\|\overline{z}^{t}\|^{2}}{3\sigma^{2}}\bigg)\bigg]$ $=1+\sum_{k=1}^{\infty}\frac{c^{k}\mathbb{E}\big[\|Y\|^{2k}\big]}{k!}\stackrel{{\scriptstyle(i)}}{{\leq}}1+\sum_{k=1}^{\infty}\frac{2dk(ck)^{k}}{k!}\stackrel{{\scriptstyle(ii)}}{{\leq}}2d\bigg(1+\frac{1}{e}\sum_{k=1}^{\infty}k(ec)^{k}\bigg),$
    where $(i)$ follows from (3), while $(ii)$ follows from $k!\geq e(k/e)^{k}$. Next, note that the sequence $\sum_{k=1}^{\infty}kp^{k}$ converges if $p<1$, with $\sum_{k=1}^{\infty}kp^{k}=\frac{p}{(1-p)^{2}}$, therefore, choosing $c<\frac{1}{e}$ in the relation above, we get
    $\mathbb{E}\bigg[\exp\bigg(\frac{cn\|\overline{z}^{t}\|^{2}}{3\sigma^{2}}\bigg)\bigg]\leq 2d\bigg(1+\frac{c}{(1-ec)^{2}}\bigg).$ (10)
    It can be readily verified that $\frac{c}{(1-ec)^{2}}\leq e-1$, for any $c\leq\frac{e-1}{e^{2}}$. Choosing $c=\frac{1}{5}$ and plugging into (10) completes the proof.
@@
where the second inequality follows from the fact that $b_{t}\in(0,1)$ and Proposition 1. Unrolling the recursion (D) and noting that $Y_{1}=X_{1}$, since $X_{1}>0$ is deterministic, we get
@@
Plugging (16) into (D) completes the proof. ∎
Appendix E Analysis Setup
@@
Appendix F Proofs for Non-convex Costs
@@
where we used Proposition 1 and (A3) in the last inequality. Plugging (23) in (F) gives the desired result. ∎
@@
Theorem 1.
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
@@
$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})+1\big)}{\sqrt{nT}}$
$+\frac{\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}\Bigg).$
@@
$=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(\frac{45\alpha_{T}^{2}\sigma^{2}L}{n}\times\frac{n\|\overline{z}^{T}\|^{2}}{15\sigma^{2}}\right)\>\Big|\>\mathcal{F}_{T}\right]\right]$
$\leq\mathbb{E}\left[\exp(B_{2,T-1})\left(\mathbb{E}\left[\exp\left(\frac{n\|\overline{z}^{T}\|^{2}}{15\sigma^{2}}\right)\>\Big|\>\mathcal{F}_{T}\right]\right)^{\nicefrac{{45\alpha_{T}^{2}\sigma^{2}L}}{{n}}}\right],$
where the last inequality follows by choosing $\alpha_{t}\leq\frac{\sqrt{n}}{3\sigma\sqrt{5L}}$ and using Proposition 1. Using the third property in Lemma 2, we get
$\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{45\alpha_{T}^{2}\sigma^{2}L}{n}\big(1+\log(2d)\big)\right)\mathbb{E}\left[\exp(B_{2,T-1})\right].$
Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{45\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}\right)$. Finally, to bound $\mathbb{E}[\exp(B_{3,T})]$, we can proceed in the same way, using the conditional independence of noise across agents (recall (A4)) to note that, if $\alpha_{t}\leq\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2/3}L^{2/3}\sqrt[3]{9}}$, then $\mathbb{E}[\exp(B_{3,T})]\leq\exp\left(\frac{9\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\right)$. Combining everything, we get
$\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg).$ (32)
@@
$\mathbb{P}\left(M_{T}>\epsilon\right)\leq\exp(-\epsilon)\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(-\epsilon+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg),$
@@
$M_{T}\leq\log(\nicefrac{{1}}{{\delta}})+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
Recalling the definition of $M_{T}$ and using the fact that the sequence of step-sizes is non-increasing we then have, with probability at least $1-\delta$
@@
$+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
@@
$+\frac{240\sigma^{2}L}{n\alpha_{T}T}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{48\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\alpha_{T}T(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
@@
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}$ $\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}$
    $+\frac{240\alpha\sigma^{2}L\big(1+\log(2d)\big)}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{15LT}}$, then with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}$ $\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}$
    $+\frac{16\sigma\sqrt{15L}\big(1+\log(2d)\big)}{\sqrt{nT}}+\frac{4n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}.$
    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{15LT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{15LT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{15L}}{\sqrt{nT}}+\frac{1}{CT}$. Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{15L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})+1\big)}{\sqrt{nT}}$
    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{4n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}.$